OpenStudy (anonymous):

Fred is standing on a 25m tower and sees a snake on the ground 18m from the base of the tower, what is the angle of depression from fred to the snake?

OpenStudy (anonymous):



OpenStudy (anonymous):

i have no idea what "angle of depression" means but to find the angle i drew, note that \[\tan(\theta)=(\frac{18}{25})=.72\] so \[\theta=\tan^{-1}(.72)\]

OpenStudy (anonymous):

if it is the other angle you need, subtract the result from 90 to get your answer

OpenStudy (saifoo.khan):

Fred should go back home and relax.

OpenStudy (mathmate):

Angle of depression is the angle it makes with the horizontal. arctan(25/18)=54.25 degrees